APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Tuesday August 17, 2010 (MJD 55426) ---=== OBSERVING TEAM ===--- Kaike Pan Audrey Simmons Paul Harding Craig Loomis French Leger (support) ---=== OBSERVING PLAN ===--- Science! ---=== OBSERVING SUMMARY ===--- Nearby thunderstorms and clouds kept us closed for the first half of the night. The last half was spent testing the new guider code. Craig and Paul worked hard on getting the 'fk5infiber' command and the calibration offsets working. Problems: The azimuth fiducials would not read during checkout, plus the rotator fiducials a little later. French was called in to fix the problem(s), but we did not catch what was wrong with them and how they were fixed, so please ask French for more details if interested. Rotator position values within the MCP periodically did not agree with it's actual position. MARVELS temperature sensor is misbehaving again. ---=== OBSERVING LOG ===--- Afternoon: ---------- 01:20Z - Started ventilating. 01:25Z - Turned amplifiers on. Checked fiducials. Azimuth fiducials don't read at all. MCP shows the position stuck at 122:18:38.68. Murmur shows: 2010-08-18 01:26:08Z sdssmcp tLatch dsc_I_trcInfo Applying correction 336738 to azimuth 2010-08-18 01:26:08Z sdssmcp tLatch dsc_E_trcErr correction for azimuth 336738 is too large Rotator fiducials also large. 01:45Z - Gathered mirror numbers. It took a long time for the secondary mirror numbers to match with the nominal values. Night: ------ Still some nearby thunderstorms, so we waited to open. 03:22Z - French both fixed the azimuth and rotator fiducials. 03:51Z - Mounted cart 17, only the guide fibers plugged. Rotator position value within the MCP does not agree with it's actual position. Tried clearing the axis to a zero value, but the position returns within a few minutes. 05:01Z - Opened enclosure. Mounted cart 12, plate 4191 for guiding tests. 05:15Z - Found guide stars on gotoField. Began guider tests with the new guider code. We called Craig for assistance. 08:50Z - Mounted cart 11, plate 4097. 11:04Z - High humidity, lots of clouds, and low telescope altitude made us close up for tonight. 11:08Z - endNight finished. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 01:47Z 62 43 6 52 (NE) -9999 - ---=== TELESCOPE STATUS ===--- 01:20Z - Began ventilation. Doors and louvers opened, fans on. 05:01Z - Opened enclosure. 11:07Z - Enclosure is on, doors and louvers are closed, and fans are off. Status at: 11:07Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 11 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. SP1 LN2 dewar scale: 258.50 lbs, 18.45 psi SP2 LN2 dewar scale: 240.80 lbs, 20.45 psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 MCP: v6_6_5 TPM: tpm_v3_5_0 STUI: 1.0b2 2010-07-21 alerts: trunk+svn114742 apo: trunk+svn114701 boss: trunk+svn115542M boss.daq: 1.2.14 boss.specMech: Offline gcamera: trunk+svn115486M guider: ** FAILED ** (running paul's version, not checked in). hub: Tag: UNTAGGED_REVISION 502:504M mcp: v6_6_5 mcp.plc: 25 platedb: trunk+svn111739 sop: v1_1_3 sos: ** FAILED ** tcc: 2.13.0 Pidlspec2d (sos3) --- v5_4_12 plate-mapper3 ------ NOCVS:v5_0_7 procedure: v3_06 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.5100 -5.5100 Axial B 16.0930 16.0940 Axial C 15.9200 15.9220 Trans D 11.0090 11.0100 Lateral E 10.6426 10.6553 Lateral F 12.2428 12.2301 GALILS Commanded: 10600. -250. 5800. 10250. 13050. 12650. Actual: 10600. -250. 5800. 10250. 13053. 12647. SETMIR VALUES PriDesOrient: 0.00 -13.67 10.53 574.80 742.37 PriOrient: 0.00 -13.67 10.53 574.80 742.37 SECONDARY: ---------- Focus: 0.00 Air Temp.: 16.6 Alt.: 30.000004 MIGS TONIGHT NOMINAL Axial A -7.8330 -7.8320 Axial B -0.0390 -0.0410 Axial C 13.8180 13.8180 Trans D 14.1810 14.1840 GALILS Commanded: 1585299. 1566704. 1598962. -1450. -50. Actual: 1585350. 1566695. 1598880. -1462. -69. SETMIR VALUES SecDesOrient: 1257.00 -9.00 -4.00 -58.00 23.67 SecOrient: 1257.01 -9.03 -3.98 -58.07 23.33 ---=== PROBLEMS IN DETAIL ===---